# On the Foundation of Distributionally Robust Reinforcement Learning

### v3

@@ 1.1 Results and Methodology
Finally, we establish counterexamples for all cases that cannot be covered by the previously mentioned principles, thus completing our unified framework. A notable insight behind the construction of these counterexamples is that the non-existence of a DPP is closely related to the controller’s ability to learn certain characteristics of the worst-case adversary over a sequence of strategically deployed actions. Therefore, an adept controller actively engages in learning, and the existence of a DPP is tied to the controller’s inability to learn within the given information structure. Consequently, in constructing counterexamples, it proves fruitful to conceptualize the controller as an agent in a multi-armed bandit or RL problem, as opposed to an MDP problem.
1.2 Criteria on the Selection of Different Formulations
In this subsection, we provide guidance on selecting the most suitable formulation from the 36 cases presented in Tables 1 to 4 for the DRRL problems. Our criteria for selection are based on four main factors: rectangularity, information structure, convexity of adversaries, and the policy classes of controllers, whether deterministic or randomized. As we shall elaborate in the proceeding discussions, a variety of attributes related to the adversary in the underlying DRMDP not only lead to a formulation that facilitates robust decision-making in a dynamic environment but also function as modeling tools to mitigate various forms of model miss-specification.
Rectangularity of the adversary:
  • SA-rectangular adversary: SA-rectangular adversary models emerge naturally when a RL agent is expect to encounter distinct states or environments after executing different actions. For instance, consider an autonomous driving agent determining whether to turn left, right, or proceed straight at an intersection. Once the agent makes a decision on the action to take, it will observe entirely different states. Therefore, it is reasonable to assume that the stochasticity in subsequent state transitions resulting from one action is unrelated of the probability structure of state transitions resulting from any other action. In such instances, SA-rectangularity could serve as an appropriate attribute of the adversary in the underlying DRRL problem. Ding et al. (2023) successfully adopted the DRRL formulation with a special SA-rectangular ambiguity set and proposed an algorithm that outperforms all baselines in the autonomous driving tasks.
  • S-rectangular adversary: Compared to SA-rectangular adversaries, S-rectangular adversaries exhibit a more constrained influence, limiting the set of transition probabilities by enforcing specific correlations among different actions. As we will demonstrate in the inventory control Example 1, in this modeling environment, it is natural to assume that at each time $t$, the adversary is restricted so that it cannot use the information of the realization today’s ordered quantity (action). This feature leads to the use of S-rectangular adversarial decision rules. Here, S-rectangularity allows the modeler to limit the power of the adversary so that it cannot use the realization of the controller’s action. Owing to this refinement, the resulting robust policies tend to be less conservative. In addition, adding reasonable constraints could facilitate the calibration of the uncertainty set; see, for example, Lotidis et al. (2023).
It is important to note that SA-rectangular adversaries are more powerful the S-rectangular counterparts, which may result in the derivation of a more conservative policy. Nevertheless, from a computational perspective, for SA-rectangular adversaries, the existence of a DPP is guaranteed (Table 1), and a $q$-function (Watkins and Dayan, 1992) with its Bellman equation is definable (equation (3.6)). These characteristics facilitate the design of model-based (Zhou et al., 2021; Panaganti and Kalathil, 2021; Yang et al., 2021; Shi and Chi, 2022; Xu et al., 2023; Shi et al., 2023; Blanchet et al., 2023) and model-free (Liu et al., 2022; Wang et al., 2023a, b; Yang et al., 2023) DRRL algorithms, enhancing the efficiency of the learning process.
Information structure of the adversary:
  • Markov time-homogeneous adversary: Markov time-homogeneous formulation is widely analyzed in the (DR)MDP literature. An adversary with this attribute is compelled to predefine one transition kernel and adhere to its transition probability across all time. Hence, such an adversary is sometimes referred to as static in the literature. Although this formulation is preferred and commonly used by practitioners and DRRL theorists due to its structural simplicity and ease of interpretation (Wiesemann et al., 2013; Zhou et al., 2021; Grand-Clement et al., 2023), we demonstrate that the corresponding DRMDP problem could pose significant computational challenges due to the absence of a DPP and, consequently, the sub-optimality of time-homogeneous policies. Specifically, we illustrate that DPPs do not exist in general if the ambiguity set is not convex (Table 3) or if the controller is deterministic (Table 4). Therefore, caution is advised for researchers and practitioners when attempting to derive optimal robust control using this formulation. In scenarios where a DPP does not exist, the development of new exact or approximation methods for computation should be attempted on a case-by-case basis.
    It’s worth noting that previous DRRL papers commonly adopt the modeling perspective of a time-homogeneous adversary (Zhou et al., 2021; Panaganti and Kalathil, 2021; Liu et al., 2022; Shi and Chi, 2022; Xu et al., 2023; Shi et al., 2023). However, these papers all refer to the optimality of deterministic Markov time homogeneous policies and Bellman equations established in Iyengar (2005), where the principal modeling assumption involves a Markov adversary and a history-dependent controller (as summarized in Table 1). In this paper, we establish the existence or non-existence of the DPP and Bellman equation for both the value and the Q-function in the case of a Markov time-homogeneous adversary, providing the theoretical foundation for these works.
  • Markov adversary: Markov adversaries is the natural modeling choice when the environment is believed to exhibit Markov transitions under the prescribed state and action spaces. However, this attribute permits non-stationary transition probabilities. Therefore, a Markov adversary could be an appropriate modeling assumption when the probability structure in the underlying Markov dynamics can vary across time. For instance, in the aforementioned inventory Example 1, the demand distribution could change from day to day, resulting in possible time-nonhomogeneity. From a computational perspective, DPPs also always exist with randomized controllers, see the second columns of Tables 1 and 3.
  • History-dependent adversary: History-dependence enables the adversary’s selection of different distributions based on various histories realized over time. Even in situations where the modeler strongly believes that the environment adheres to Markov dynamics, the presence of partial observation or incomplete modeling of the system state usually results in the optimal controller being history-dependent. This, in turn, gives rise to non-Markov state transition dynamics. Therefore, formulating the DRRL problem with a history-dependent adversary can not only fortify the optimal controller against potential non-Markov shifts in the transition structures of the environment, but also can be wielded as a tool to hedge against errors stemming from the use of a Markov model in cases where the underlying Markov states are misspecified or partially observable.
    Moreover, this attribute of the adversary is also relevant in environments where the underlying dynamics or the data generating process is not an MDP, but the practitioner chooses to employ a DRRL model as an approximation due to computational or data efficiency considerations. This situation arises, for instance, when the RL agent attempts to achieve learning and optimize its policy through strategic interaction with the environment. For example, consider an RL agent learning bidding strategies in online advertising auctions (Cai et al., 2017). In recent years, the industry has witnessed a paradigm shift, transitioning from the established second-price auctions to the first-price auction model [^1] https://blog.google/products/adsense/our-move-to-a-first-price-auction/. In second-price auctions, it is widely recognized that truthful bidding is a weakly dominant strategy; however, this is not the case in first-price auctions. Suppose that the RL agent learns from an exchange with second-price auctions and is poised to deploy a policy in another exchange governed by first-price auctions. In this scenario, the environment comprises numerous competing strategic users, potentially reacting distinctively to the RL agent’s varied actions. As the competing strategic users are learning to respond to the strategy of the agent in the first-price auctions, they may take into account the historical behavior of the RL agent. Therefore, this may render the underlying environment non-Markov. In this setting, a non-Markov control formulation that specifically models this learning process might be excessively intricate to formulate, learn, or optimize, while a DRRL formulation that exhibits a DPP could serve as a computationally feasible approximation.
    Furthermore, from a computational standpoint, this formulation is also advantageous, as evidenced by the guaranteed existence of DPPs, highlighted in the first columns of Tables 1 to 4.
We remark that, by symmetry, the controller can also adopt these attributes pertaining their information availability. The rationale behind selecting these information structures for the controller mirrors that for the adversary. Consequently, we opt not to delve into a separate discussion on this matter.
Convexity of the adversarial ambiguity sets:
  • Convex adversary: DRMDP and DRRL models with convex adversaries are prevalent in literature, with instances spanning various ambiguity sets such as optimal-transport-based (Taşkesen et al., 2023; Yang, 2021), convex $f$-divergence-based (Yang et al., 2021; Shi et al., 2023), TV-distance-based (Shi et al., 2023), as well as the special ambiguity set employed in Xu and Mannor (2010); Wiesemann et al. (2013); Ding et al. (2023). Our findings, as presented in Tables 1 and 2, further underscore the computational benefits of convex adversaries. Nevertheless, a word of caution is necessary when navigating scenarios involving a deterministic controller paired with an S-rectangular convex adversary, as the DPP may not be applicable in this case.
  • Non-convex adversary. A non-convex adversary could be a modeling choice when the underlying system dynamics is believed to be deterministic (Post and Ye, 2013). Also, the utilization of a non-convex adversary can also stems from the adoption of unions of convex adversarial ambiguity sets. This choice may be the consequence of a data-driven calibration aiming to adjust the size and geometry of ambiguity sets, as proposed by Hong et al. (2020). From a computational standpoint, the presence of non-convex adversaries not only introduces complexity in solving the distributionally robust variant of the Bellman equations for both the value (3.1) and the $q$-function (3.4), but our research findings also suggest that the DPP may not be applicable at all. This observation holds true for both deterministic and randomized controller policies, as summarized in Table 3 and Table 4.
Possibility of randomization (convexity) of the controller:
  • The randomized controller policy class: While it is well-known that deterministic policies could achieve optimality in the traditional (non-DR) MDP setting, it is in general not the case under the presence of a S-rectangular adversary. This behavior is first observed and analyzed in Wiesemann et al. (2013). Hence, in the pursuit of optimal robust control or DRRL, the freedom to choose randomized controller policies becomes essential. Furthermore, from the perspective of the existence of a DPP, randomized controllers offer an additional guarantee, when a history-dependent controller policies is paired with Markov adversaries. This is evident from a comparison of Table 3 and 4.
  • The deterministic controller policy class: While optimality may not always be achieved with deterministic policies, implementing randomized policies can prove challenging in various real-world scenarios. Take, for instance, the context of medical treatment planning, where a patient’s current health condition dictates the selection from a range of dynamic treatment regimes (Chakraborty and Murphy, 2014). In such a scenario, a randomized policy would necessitate assigning probability distributions across various treatments for a specific health state, a practice that is neither legally permissible nor ethically acceptable.
    However, when restricting our focus to deterministic policies, it is important to note that the absence of a DPP implies that Markov time-homogeneous policies are not assured to be optimal in numerous scenarios, even in the presence of a convex adversary. This is highlighted in Table 4. To pursue higher rewards, the controller might explore the implementation of non-stationary or even history-dependent policies.
It’s worth noting that as pointed out in the previous section, while we currently introduce only randomized and deterministic controller policies, the subsequent sections will build on the convexity of the controller policy classes. Specifically, the randomized controller policy class represents a convex controller, whereas the deterministic controller policy class is an instance of a non-convex controller. The findings presented in Table 3 and 4 extend to general convex and non-convex controller policy classes, respectively. From both a modeling and learning standpoint, the presence of a DPP for a set of controller policies with general geometries opens avenues for potential extensions of DRRL algorithms. These extensions can be tailored to scenarios where the controller is constrained to employ only a subset of randomized decisions.
@@ 1.3 Literature Review
This paper contributes to the existing literature by offering a unified perspective on the existence of a DPP in the context of DRMDP, encompassing various information structures and adversarial rectangularity. This unification leads to the identification of new results regarding the existence and non-existence of DPP under various policy structures of the controller and the adversary. To appreciate this contribution, we begin with a concise overview of the relevant literature in this domain.
@@
The exploration of DRMDP models is motivated by the pursuit of optimal decision-making in stochastic environments, where the underlying probability structure is shaped by a decision maker (referred to as the controller) and an opposing counterpart (referred to as the adversary), engaged in a zero-sum interaction. This naturally aligns with the literature on two-player zero-sum stochastic games (Shapley, 1953; Solan and Vieille, 2015), a connection we will explore subsequently in Section 6.2. However, unlike traditional stochastic games, in DRMDP, the adversary’s choices are confined to a predetermined subset of transition distributions, while the controller typically still has the liberty to choose an arbitrary randomized action.
The presence of a DPP is a highly desirable property of the (DR)MDP. With the DPP in place, there exists an optimal robust policy for the controller that is stationary Markov deterministic. The DPP not only offers valuable theoretical insights into the underlying dynamics but also serves as a crucial computational tool for achieving policy learning, with or without full knowledge of the probability structure of the underlying model. Therefore, the focus of DRMDP studies is on establishing conditions that facilitate the presence of a DPP while maintaining the necessary structural integrity tailored to the modeled environments. To achieve this, previous works considered adversarial environments exhibiting a property known as rectangularity. The concept of a rectangular adversary was first introduced by Iyengar (2005) and Nilim and El Ghaoui (2005). In these foundational works, what is now recognized as an SA-rectangular adversary was considered, and the existence of a DPP was demonstrated, as summarized in Table 1. This assumption also underlies the pioneering work in min-max control by González-Trejo et al. (2002), where the same DPP is established within the context of a history-dependent adversary.
@@
However, as illustrated in Example 1 below, the SA-rectangular adversary may wield excessive power in many practical modeling environments. This leads to the introduction of S-rectangularity by Le Tallec (2007), Xu and Mannor (2010), and Wiesemann et al. (2013). Notably, these works consider specific convex S-rectangular adversaries and establish the existence of a DPP, as outlined in Table 2. Unlike the SA-rectangular setting, Wiesemann et al. (2013) shows that, in general, the optimal stationary Markov robust policy needs to be randomized.
@@
Building on this, Le Tallec (2007) and Wiesemann et al. (2013) also introduce the concept of a general rectangular adversary and establish the non-existence of a canonical DPP. To further constrain the adversary’s power and preserve the structural integrity of the modeled environment, the recent work by Goyal and Grand-Clément (2023) introduces a new notion known as R-rectangularity.
@@
In addition to rectangularity, another crucial factor influencing the existence or non-existence of the DPP is the asymmetry in information availability between the controller and the adversary. Iyengar (2005) noted, without offering a proof, that the asymmetry in historical information structure between the controller and the adversary dynamics could lead to the non-existence of a stationary optimal policy for maximizing infinite-horizon discounted rewards. This insight is also noted by Wiesemann et al. (2013, Table 1, S-rectangular, nonconvex) without giving a counterexample. The work by González-Trejo et al. (2002) approaches the min-max control problem with symmetric history-dependent controllers and adversaries. On the other hand, Iyengar (2005); Nilim and El Ghaoui (2005); Xu and Mannor (2010); Wiesemann et al. (2013) take into account non-symmetric information structures, as indicated in the tables.
Much like the formulations of DRMDPs using the language of classical MDPs, there exists a closely related line of research to DRMDP that studies the distributionally robust variant of multistage stochastic programs (MSPs), see for example Huang et al. (2017); Shapiro (2021); Pichler and Shapiro (2021), Shapiro et al. (2021, Chapter 7) and references therein. While this literature primarily focuses on finite-horizon environments, it shares strong connections with our DRMDP framework. In this paper, we provide a discussion on the reformulation of our DRMDP problems with history-dependent controllers and adversaries as distributionally robust MSPs.
@@
There exists a substantial body of literature dedicated to investigating the finite-time sample complexity of achieving optimality in tabular DRRL environments. This line of research crucially relies on DPPs and specifically focuses on divergence-based ambiguity sets of SA-rectangular adversaries. In line with the classical tabular RL setting, the approaches in these algorithms can be categorized based on whether the algorithm explicitly estimates the entire transition kernel. If it does, it falls under the scope of a model-based approach. Notable papers adopting a model-based approach include Zhou et al. (2021); Panaganti and Kalathil (2021); Yang et al. (2021); Shi and Chi (2022); Xu et al. (2023); Shi et al. (2023); Blanchet et al. (2023). Conversely, the literature exploring model-free algorithms includes works such as Liu et al. (2022); Wang et al. (2023a, b); Yang et al. (2023). In the specific domain of health care, Saghafian (2023) proposes RL methods to learn an optimal dynamic treatment regime in an ambiguous environment.
@@ 1.4 Remark on Paper Organization
The structure of the paper unfolds as follows: Section 2 formulates DRMDP problems, encompassing various types of controllers and adversaries. Our key findings for scenarios where the DPP holds with full generality are expounded in Section 3. Section 4 delves into the implications of the DPP, specifically addressing the optimality of Markov time-homogeneous policies. In Section 5, we provide counterexamples for cases where the DPP does not hold. Sections 6.1 and 6.2 explore the connections between our DRMDP formulations and distributionally robust multistage stochastic programs and stochastic games, respectively. Finally, in Section 7, we draw conclusions and outline potential avenues for future research.
2 DRMDPs: Construction and Definitions
While the theory we present in this paper naturally extends to infinite (countable, continuum, or general measurable) state and action spaces, to facilitate a better understanding of the readers from multiple disciplines, we focus our discussion exclusively on the setting of finite state and action spaces. In particular, let $S$ and $A$ denote the finite state and action spaces, respectively, and equip them with $\sigma$-fields $\mathcal{S}=2^{S}$ and $\mathcal{A}=2^{A}$. It should be noted that the formulation in this section is valid for abstract state and action measurable spaces $(S,\mathcal{S})$ and $(A,\mathcal{A})$.
In this context of finite states and actions, we consider the canonical space of an MDP. Specifically, let $(\Omega=(|S|\times|A|)^{\mathbb{Z}_{\geq 0}},\mathcal{F})$ be the underlying measurable space where $\mathcal{F}$ is the cylinder $\sigma$-field. Define the stochastic process $\left\{{(X_{t},A_{t}),{t\geq 0}}\right\}$ by the point evaluation $X_{t}(\omega)=s_{t},A_{t}(\omega)=a_{t}$ for all $t\geq 0$ and any $\omega=(s_{0},a_{0},s_{1},a_{1},\dots)\in\Omega$. In classical MDP formulations, the policy of the controller induces a probability measure on $(\Omega,\mathcal{F})$. In comparison, within the realm of robust MDPs, the interplay between the controller and the adversary dictates a measure on the sample path space. Consequently, to rigorously formulate a robust MDP, we commence by introducing the policy classes of both the controller and the adversary.
Similar to regular MDP formulations, a controller typically has the flexibility to choose its action at the current time and space based on all historical information (i.e. the state-action sequence realized until the current time instance) available. In a robust MDP, both the controller and the adversary, in general, are empowered to employ history-dependent policies as well. To establish a formulation of robust MDP with this characteristic, we introduce the definitions of the controller’s and adversary’s history and their respective notions of history-dependence in the context of a robust MDP setting.

### v4

@@ 1.1 Results and Methodology
Moreover, we establish counterexamples for all cases that cannot be covered by the previously mentioned principles, thus completing our unified framework. A notable insight behind the construction of these counterexamples is that the non-existence of a DPP is closely related to the controller’s ability to learn certain characteristics of the worst-case adversary over a sequence of strategically deployed actions. Therefore, an adept controller actively engages in learning, and the existence of a DPP is tied to the controller’s inability to learn within the given information structure. Consequently, in constructing counterexamples, it proves fruitful to conceptualize the controller as an agent in a multi-armed bandit or RL problem, as opposed to an MDP problem.
These counterexamples highlight the fundamental limitations of non-history-adaptive policies in RMDPs where Markovian or time-homogeneous adversaries select actions from non-convex S-rectangular sets. A particularly relevant setting of practical interests involves history-dependent controllers—either randomized or deterministic—interacting with time-homogeneous adversaries. To address this challenge, we propose an RL-based history-dependent policy that is both provably efficient and practically implementable (c.f. Theorem 6). Our policy achieves asymptotic optimality as the effective horizon $1/(1-\gamma)$ grows and builds on model-based RL algorithms that are widely used in practice. Indeed, it exploits the adversary’s time-homogeneity to enable strategic adaptation, ensuring sufficient exploration and the near-optimal use of historical data.
@@ 1.2 Literature Review
This paper contributes to the existing literature by offering a unified perspective on the existence of a DPP in the context of RMDP, encompassing various information structures and adversarial rectangularity. This unification leads to the identification of new results regarding the existence and non-existence of DPP under various policy structures of the controller and the adversary. To appreciate this contribution, we begin with a concise overview of the relevant literature in this domain.
@@
RMDP models are motivated by optimal decision-making in stochastic environments where the underlying probability structure is shaped by a decision maker (referred to as the controller) and an opposing counterpart (referred to as the adversary), engaged in a zero-sum interaction. This naturally aligns with the literature on two-player zero-sum stochastic games [64, 71], a connection we will explore subsequently in Section 7.2. However, unlike traditional stochastic games, in RMDP, the adversary’s choices are confined to a predetermined subset of transition distributions, while the controller typically still has the liberty to choose an arbitrary randomized action.
In RMDPs, DPP is highly desirable because it certifies the optimality of stationary Markov policies and enables efficient control and RL algorithms. Accordingly, research has focused on conditions that retain realistic structural features yet still guarantee a DPP—most notably by imposing the adversary’s rectangularity property. This concept was first introduced by Iyengar, [28] and Nilim and El Ghaoui, [47]. In these foundational works, what is now recognized as an SA-rectangular adversary was considered, and the existence of a DPP was demonstrated, as summarized in Table 1. This assumption also underlies González-Trejo et al., [18] on min-max control, where the same DPP is established within the context of a history-dependent adversary.
@@
However, as illustrated in Example 1 below, the SA-rectangular adversary may wield excessive power in many practical modeling environments. This leads to the introduction of S-rectangularity [34, 80, 79]. Notably, these works consider specific convex S-rectangular adversaries and establish the existence of a DPP, as outlined in Table 2. Unlike the SA-rectangular setting, Wiesemann et al., [79] shows that, in general, the optimal stationary Markov robust policy needs to be randomized.
@@
Building on this, Le Tallec, [34] and Wiesemann et al., [79] also introduce the concept of a general rectangular adversary and establish the non-existence of a canonical DPP. To further constrain the adversary’s power and preserve the structural integrity of the modeled environment, the recent work by Mannor et al., [41], Goyal and Grand-Clément, [19] introduces a new notion known as k- and r-rectangularity.
@@
In addition to rectangularity, another crucial factor influencing the existence or non-existence of the DPP is the asymmetry in information availability between the controller and the adversary. Iyengar, [28] noted, without offering a proof, that the asymmetry in historical information structure between the controller and the adversary dynamics could lead to the non-existence of a stationary optimal policy for maximizing infinite-horizon discounted rewards. This insight is also noted by Wiesemann et al., [79, Table 1, S-rectangular, nonconvex] without giving a counterexample. The work by González-Trejo et al., [18] approaches the min-max control problem with symmetric history-dependent controllers and adversaries. On the other hand, Iyengar, [28], Nilim and El Ghaoui, [47], Xu and Mannor, [80], Wiesemann et al., [79] take into account non-symmetric information structures, as indicated in the tables.
@@
There exists a substantial body of literature dedicated to investigating the finite-time sample complexity of achieving optimality in tabular DRRL environments. This line of research crucially relies on DPPs and specifically focuses on divergence-based ambiguity sets of SA-rectangular adversaries. In line with the classical tabular RL setting, the approaches in these algorithms can be categorized based on whether the algorithm explicitly estimates the entire transition kernel. If it does, it falls under the scope of a model-based approach. There is a rapidly growing literature, adopting either a model-based [87, 51, 84, 65, 81, 66, 5] or a model-free [39, 76, 77, 83] approach. In the specific domain of health care, Saghafian, [57] proposes RL methods to learn an optimal dynamic treatment regime in an ambiguous environment.
Much like the formulations of RMDPs using the language of classical MDPs, there exists a closely related line of research to RMDP that studies the robust variant of multistage stochastic programs (MSPs), see for example [27, 59, 53, 61] and references therein. While this literature primarily focuses on finite-horizon environments, it shares strong connections with our RMDP framework and can be seen as alternative interpretations for DRRL models.
Moreover, robust MSPs are closely related to risk measures. In the static setting, certain distributionally robust objectives are known to correspond to coherent risk measures. A well-known example is that the conditional value-at-risk (CVaR) measure is equivalent to a distributionally robust objective with the ambiguity set $\mathcal{P}=\left\{Q:dQ/dP\in[0,1/(1-\alpha)],Q(\Omega)=1\right\}$ In the dynamic setting, Le Tallec, [34], Osogami, [49], Pichler and Shapiro, [53], Shapiro and Pichler, [63] show that robust multistage programming is equivalent to optimization with respect to nested (also called iterated or dynamic) conditional risk functionals [56].
Another related line of research is risk-sensitive Markov Decision Processes (MDPs). In risk-sensitive MDPs, the objective is to optimize $E\left[U\left(\sum_{k=0}^{\infty}\gamma^{k}r(X_{t},A_{t})\right)\right],$ where $U$ is a nonlinear utility function [24, 16, 12, 13, 2]. A common choice would be $U(\cdot)=\exp(\cdot)$. The risk-aware MDP literature also study objectives of the form $\rho\left(\sum_{k=0}^{\infty}\gamma^{k}r(X_{t},A_{t})\right),$ where $\rho$ is a risk measure [23, 73, 62]. Notably, Chow et al., [9] showed that risk-sensitive MDPs with CVaR risk measure is equivalent to robust MDPs with multiplicative probability perturbation. In addition, Delage and Mannor, [11], Haskell and Jain, [22], Chow et al., [8] studied the risk-constrained MDPs, where they maximize expected cumulative rewards under risk measure constraints.
Establishing a dynamic programming principle in risk-sensitive MDPs often requires augmenting the state space to maintain the necessary Markov property, which differs from the setting in this paper. Moreover, unlike robust MDPs, risk-sensitive MDPs do not involve ambiguity sets or adversarial agents. As a result, there is no issue of information asymmetry in this framework.
@@ 1.3 Remark on Paper Organization
The structure of the paper unfolds as follows: Section 2 formulates RMDP problems, encompassing various types of controllers and adversaries. Our key findings for scenarios where the DPP holds with full generality are expounded in Section 3. Section 4 delves into the implications of the DPP, specifically addressing the optimality of Markov time-homogeneous policies. In Section 5, we provide counterexamples for cases where the DPP does not hold. Section 6 proposes and analyzes an efficient RL-based history-dependent policy that achieves asymptotically optimal performance against a time-homogeneous non-convex adversary. Sections 7.1 and 7.2 explore the connections between our RMDP formulations and robust multistage stochastic programs and stochastic games, respectively. Finally, in Section 8, we offer insights to guide RMDP model selection for practitioners and outline several promising directions for future research.
2 Robust MDPs: Construction and Definitions
While the theory we present in this paper naturally extends to infinite (countable, continuum, or general measurable) state and action spaces, to facilitate a better understanding of the readers from multiple disciplines, we focus our discussion exclusively on the setting of finite state and action spaces. In particular, let $S$ and $A$ denote the finite state and action spaces, respectively, and equip them with the $\sigma$-fields of all subsets $\mathcal{S}=2^{S}$ and $\mathcal{A}=2^{A}$. It should be noted that the formulation in this section is valid for abstract state and action measurable spaces $(S,\mathcal{S})$ and $(A,\mathcal{A})$.
In this context of finite states and actions, we consider the canonical space of an MDP. Specifically, let $(\Omega=(|S|\times|A|)^{\mathbb{Z}_{\geq 0}},\mathcal{F})$ be the underlying measurable space where $\mathcal{F}$ is the cylinder $\sigma$-field. Define the stochastic process $\left\{{(X_{t},A_{t}),{t\geq 0}}\right\}$ by the point evaluation $X_{t}(\omega)=s_{t},A_{t}(\omega)=a_{t}$ for all $t\geq 0$ and any $\omega=(s_{0},a_{0},s_{1},a_{1},\dots)\in\Omega$. In classical MDP formulations, the policy of the controller induces a probability measure on $(\Omega,\mathcal{F})$. In comparison, within the realm of RMDPs, the interplay between the controller and the adversary dictates a measure on the sample path space. Consequently, to rigorously formulate an RMDP, we commence by introducing the policy classes of both the controller and the adversary.
Similar to regular MDP formulations, a controller typically has the flexibility to choose its action at the current time and space based on all historical information (i.e. the state-action sequence realized until the current time instance) available. In an RMDP, both the controller and the adversary, in general, are empowered to employ history-dependent policies as well. To establish a formulation of RMDP with this characteristic, we introduce the definitions of the controller’s and adversary’s history and their respective notions of history-dependence in the context of an RMDP setting.